Ticket Number: HEA-NOC/20080122-9 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: EPA
Ticket Opened: 20080122 15:06 UTC Problem Start: 20080122 15:00 UTC
Ticket Update: 20080124 15:32 UTC Problem End: 20080122 18:00 UTC
EPA are testing a new firewall
20080122 15:05 COC Monitoring. EPA will inform us when they have completed
20080122 16:07 COC EPA requested logs and configuration details of their
Sent logs as requested to Entropy.
20080124 15:29 COC EPA are still having connectivity issues between their
Firewall between Dublin and Wexford. Connectivity is still Intermittent.
This ticket can be monitored at http://www.hea.net/tickets/20080122-9
HEAnet Limited. Registered in Ireland, No. 275301.